# Autonomation, not Automation: Activities and Needs of Fact-checkers as a Basis for Designing Human-Centered AI Systems

### v2

@@ Abstract.
To mitigate the negative effects of false information more effectively, the development of Artificial Intelligence (AI) systems assisting fact-checkers is needed. Nevertheless, the lack of focus on the needs of these stakeholders results in their limited acceptance and skepticism toward automating the whole fact-checking process. In this study, we conducted semi-structured in-depth interviews with Central European fact-checkers. Their activities and problems were analyzed using iterative content analysis. The most significant problems were validated with a survey of European fact-checkers, in which we collected 24 responses from 20 countries, i.e., 62% of active European signatories of the International Fact-Checking Network (IFCN).
Our contributions include an in-depth examination of the variability of fact-checking work in non-English speaking regions, which still remained largely uncovered. By aligning them with the knowledge from prior studies, we created conceptual models that help understand the fact-checking processes. Thanks to the interdisciplinary collaboration, we extend the fact-checking process in AI research by three additional stages. In addition, we mapped our findings on the fact-checkers’ activities and needs to the relevant tasks for AI research. The new opportunities identified for AI researchers and developers have implications for the focus of AI research in this domain.
@@
ccs: Human-centered computingccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Research shows that the most effective mitigation of political misinformation in terms of lowering issue agreement and perceived accuracy is a combination of information literacy interventions and fact-checkers’ efforts (Hameleers, 2022). Fact-checkers have an irreplaceable role in mitigating the effects of misinformation, as they successfully lower the agreement with attitudinal congruent political misinformation and can help overcome political polarization (Hameleers and van der Meer, 2020).
@@
A fact-checker is “a person whose job is to make sure that the facts are correct, especially in something published”  (noa, [n. d.]). These professionals work either in larger newspaper agencies (e.g., in AFP, Deutsche Welle, Washington Post) or small or medium-sized NGOs focused just on fact-checking (e.g., Full Fact, PolitiFact.com, FactCheck.org, etc.). Many aspects of the fact-checker’s work, as well as the problems this profession faces differ across the positions (e.g., the required education), remain unclear (e.g., stages in the fact-checking process, see Fig. 3), or have not been discovered at all (e.g., geo-cultural differences – see Fig. 1).
@@
Although growing, the number of professional fact-checkers remains low in comparison to the vast amount of misinformation. The latest census by the Duke Reporters’ Lab identified 378 active fact-checking projects worldwide. Only 88 organizations worldwide (34 in Europe) actively cooperate with social media platforms through the International Fact-checking network and in some cases, there is merely one fact-checker per country. The imbalance between the scarcity of fact-checkers and misinformation overload causes that just a fraction of potential false content is being checked. Furthermore, the work of fact-checkers is laborious and sometimes repetitive. There is, however, an opportunity to improve the balance by providing appropriate technological support for challenging or repetitive tasks.
@@
While there is an increasing body of research on automation of the fact-checking process, especially with the utilization of Artificial Intelligence (AI) technologies, full automation of the whole fact-checking process is still viewed with skepticism by the fact-checkers as well as researchers (noa, 2020; Micallef et al., 2022; Dierickx, 2022; Juneja and Mitra, 2022). The main reason is that some parts of the fact-checking process require human judgments or actions. Tasks such as assessing the credibility of the sources of evidence are currently not fully automated with sufficient trustworthiness (e.g., by providing appropriate explanations), accuracy, or generability (e.g., checking complex recent claims where evidence may still be missing) (Nakov et al., 2021a). The real potential of AI-based tools, therefore, currently lies in assisted fact-checking instead of pursuing to automate the whole process.
@@
With this in mind, we reinvoke a philosophy of machine design called “autonomation”, developed in the Toyota Production System (Ohno and Bodek, 1988). Autonomation is a blend of “autonomous” and “automation” and may be described as “intelligent automation” or “automation with a human touch”. Compared to (semi-)automation that is typically orchestrated by a centralized computer controller, autonomation separates the work of humans from machines, and thus enables the quick correction of the mistakes made by machines. Yet, autonomation relieves humans of the need to continuously judge whether the operation of the machine is right. In the “autonomated” systems, the workers are self-inspecting their work and can source-inspect the work of the machines. This is a difference from an “automatic” system (that operates without human intervention) or an “automated” system (that requires human input or monitoring but is controlled by technology). Autonomation liberates people from automatable tasks, whereas technology has to serve people and processes, not vice versa.
@@
Although, the HCI community has significantly helped to better understand the gap between the social and technical, the existing research works on AI-assisted fact-checking are still often detached from real fact-checkers, and thus do not optimally comply with their actual needs and expectations. As Nakov et al. point out: there is a “lack of collaboration between researchers and practitioners in defining tasks and developing datasets” for automated fact-checking (Nakov et al., 2021a). This contributes to the skepticism of fact-checkers towards automating the entire fact-checking process, especially processes that require human judgments (noa, 2020; Micallef et al., 2022). On the contrary, some studies, e.g. (Nguyen et al., 2018) mention too much trust in the system’s predictions.
@@
In this study, we aspire to research the fact-checking procedures to design the effective and appropriate human-centered AI tools for countering false information. The role of humans is central here: the tools should empower humans instead of replacing them. This work offers the following contributions:
@@
    We unify the various categorizations of the fact-checking process across different computer and social science literature  (noa, 2020; Nakov et al., 2021a; Micallef et al., 2022; Barrón-Cedeño et al., 2020) and we extend the process in AI research by three additional stages.
@@
    We investigate the activities, problems, resources and tools of under-researched European fact-checkers and examine the specifics of fact-checking work in the non-English speaking regions. We align them with the knowledge from existing studies, namely (noa, 2020; Nakov et al., 2021a; Dierickx, 2022; Micallef et al., 2022), and visualize them jointly as conceptual models.
@@
    We identify new opportunities for research and development of AI tools that would be useful in the day-to-day work of fact-checkers following the principles of human-centered AI.
2. Related work
A lot of AI research is devoted to assisting fact-checkers by automating the individual stages of fact-checkers’ work (Nakov et al., 2021a; Zeng et al., 2021). It focuses on tasks such as finding claims worth fact-checking (Gencheva et al., 2017; Berendt et al., 2021; Konstantinovskiy et al., 2021), searching for previously fact-checked claims (Shaar et al., 2021a; Vo and Lee, 2020; Sheng et al., 2021; Kazemi et al., 2021; Hardalov et al., 2022), or claim verification and evidence retrieval (Zhao et al., 2020; Lee et al., 2020; Popat et al., 2017). Another group of research focuses on the development of the necessary datasets (Shaar et al., 2020; Nakov et al., 2021b; Srba et al., 2022; Hardalov et al., 2022) and end-to-end systems or monitoring platforms (Hassan et al., 2017; Srba, [n. d.]). The NLP (natural language processing) technology capabilities and limitations for human-centered automation in fact-checking were also summarized (Das et al., 2023). Developing AI-based methods has also been addressed through competitions and data challenges, especially as a part of CheckThat! Labs, where tasks such as check-worthiness estimation (Shaar et al., 2021b) or detecting previously fact-checked claims (Shaar et al., 2021a) have been proposed.
Some research is also devoted to the routines and challenges of fact-checkers. The Full Fact report (noa, 2020) was the first one that globally examined the practices of today’s fact-checkers. Following the semi-structured interviews with fact-checkers from 19 organizations, the report lays out the main challenges fact-checkers face, such as large amounts of potential claims to check. The report was complemented by (Nakov et al., 2021a) with an analysis of current technological capabilities for fact-checking automation. The recognized potential of AI-based solutions lies in multiple areas: 1) finding claims from the large information ecosystem; 2) finding previously fact-checked claims; 3) finding supporting evidence (in a form of text, audio or video), translating (for multilingual content) and summarizing relevant posts, articles, and documents if needed; and finally, 4) detecting claims that are spreading fast to slow them down.
Further research on fact-checkers’ practices was done by NORDIS (NORdic observatory for digital media and information DISorders) and focused on the tools for some parts of the fact-checking process (Dierickx, 2022). Semi-structured interviews with 14 respondents from Nordic countries (Norway, Sweden, Finland, Denmark), supplemented with 5 non-Nordic professional journalists and fact-checkers, were conducted. Another semi-structured interview with 21 fact-checkers from 19 countries (Micallef et al., 2022) points to the motivations of fact-checkers and the reasons for the limited uptake of computational tools by these professionals. Finally, the research of (Juneja and Mitra, 2022) investigates the fact-checking infrastructures and thus expands the perception of the current fact-checkers’ work.
Although the research of fact-checkers’ practices and problems is growing, the European region as well as non-American and non-English speaking fact-checkers remain still under-explored as can be seen on the map (Fig. 1). Secondly, the aforementioned studies did not aspire to investigate the possible geographical, language, or organizational differences in the problems of fact-checkers. Thirdly, despite existing analyses of fact-checkers’ work, there is a partial disconnection between the fact-checkers’ needs and the efforts to create (AI) tools to support their work. Their design is treated as a technical solution to a technological problem and ignores the complex social and situational context  (Juneja and Mitra, 2022). There is also a lack of collaboration between AI researchers and developers with fact-checkers (Nakov et al., 2021a). For example, the stage monitoring the online space is insufficiently covered by AI research in the context of fact-checkers, although carried out routinely by these professionals and in some works (noa, 2020; Juneja and Mitra, 2022) even recognized as “the hardest part of the fact-checking process”.
@@ 2. Related work
Figure 1. The regions covered by available research of fact-checkers
@@
This problem is not common only to fact-checkers. Although the conceptual foundations of HCAI are extensively discussed in recent literature  (Shneiderman, 2022; Xu, 2019) and guidelines for building human-centered AI products exist (e.g. (Google, 2021)), the industry practices and methods appear to lag behind (Hartikainen et al., 2022; Bingley et al., 2023).

### v3

# Autonomation, Not Automation: Activities and Needs of European Fact-checkers as a Basis for Designing Human-Centered AI Systems

@@ Abstract.
To mitigate the negative effects of false information more effectively, the development of Artificial Intelligence (AI) systems to assist fact-checkers is needed. Nevertheless, the lack of focus on the needs of these stakeholders results in their limited acceptance and skepticism toward automating the whole fact-checking process. In this study, we conducted semi-structured in-depth interviews with Central European fact-checkers. Their activities and problems were analyzed using iterative content analysis. The most significant problems were validated with a survey of European fact-checkers, in which we collected 24 responses from 20 countries, i.e., 62% of active European signatories of the International Fact-Checking Network (IFCN). Our contributions include an in-depth examination of the variability of fact-checking work in non-English-speaking regions, which still remained largely uncovered. By aligning them with the knowledge from prior studies, we created conceptual models that help to understand the fact-checking processes. In addition, we mapped our findings on the fact-checkers’ activities and needs to the relevant tasks for AI research, while providing a discussion on three AI tasks that were not covered by previous similar studies. The new opportunities identified for AI researchers and developers have implications for the focus of AI research in this domain.
@@
copyright: acmlicensedjournal: JRCjournalyear: 2025journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/3764592ccs: Human-centered computingccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
A fact-checker is “a person whose job is to make sure that the facts are correct, especially in something published”  (noa, [n. d.]). These professionals work either in larger newspaper agencies (e.g., in AFP, Deutsche Welle, Washington Post) or small or medium-sized NGOs focused just on fact-checking (e.g., Full Fact, PolitiFact.com, FactCheck.org, etc.). Many aspects of the fact-checker’s work, as well as the problems this profession faces, differ across the positions (e.g., the required education) or remain unclear (e.g., stages in the fact-checking process; see Fig. 3 and geo-cultural differences – see Fig. 1). Nevertheless, research shows (Koliska and Roberts, 2024) that fact-checkers define their role and work as a public service and develop activities to instill a culture of factuality that they value. Exemptions include fringe state‐supported fact‐checkers associated with authoritarian governance structures that weaponize fact‐checking practices and exploit or mimic the social standing of accredited fact‐checkers to fact‐check for national interests (Montaña-Niño et al., 2024).
Research literature reports varying findings on the extent to which fact-checking reduces misinformation, with estimates ranging from minor effects (Clayton et al., 2020), to moderate (Porter and Wood, 2021), to substantial (Boukes and Hameleers, 2023) reductions in belief in misinformation after exposure to a fact-checking label. In these studies, participants assessed the accuracy of various real and fabricated statements, either with or without a fact-checking tag. Such discrepancies may stem from contextual differences in how misinformation is evaluated, including the type of misinformation, the platform and format of the fact-check, the source of the debunking information, and the metrics used to measure effectiveness (ranging from a 3-point Likert scale on the falsity of the news (Clayton et al., 2020) to a 7-point Likert scale on agreement with the news  (Boukes and Hameleers, 2023)), as well as prior beliefs or personality characteristics of the participants. For example, knowledgeable individuals with partisan views tend to scrutinize fact-checking more  (Walter et al., 2020). Also, who the fact-checker is matters when assessing the success in debunking misperceptions (Van Erkel et al., 2024). However, in general, fact-checking has been shown to be effective in combination with information literacy interventions in reducing political misinformation in terms of lowering issue agreement and perceived accuracy  (Hameleers, 2022). Fact-checkers also successfully lowered the agreement with attitudinally congruent political misinformation and can help overcome political polarization (Hameleers and van der Meer, 2020).
Despite these positive outcomes of fact-checking, active cooperation with social media platforms such as X or Facebook has been ceased recently or is in danger of being ceased despite previous statements that ‘it works’ [^1] https://web.archive.org/web/20250110130651/https://www.facebook.com/formedia/blog/third-party-fact-checking-how-it-works. Without the cooperation with social media, the dissemination of fact-checks will be even less effective. Also, there is currently a trend of replacing the professional fact-checks by community notes, i.e., the consensus of regular users. On one hand, community notes might be an effective approach to mitigate trust issues with simple misinformation flags (Drolsbach et al., 2024). On the other hand, it may result in the fact-checking process being more subjective and susceptible to the bias of community members. More specifically, previous studies showed that community notes tend to be biased towards fact-checking posts from large accounts (Pilarski et al., 2024). Also, contextual features – in particular, the partisanship of the users – are far more predictive of judgments than the content of the posts and evaluations themselves (e.g., users preferentially challenge content from those with whom they disagree politically) (Allen et al., 2022). Finally, a recent study (Borenstein et al., 2025) showed that community notes are more of a complement than a replacement for traditional fact-checking since community moderation relies on professional fact-checking – more specifically, community notes on posts linked to broader narratives are twice as likely to reference fact-checking sources compared to other sources.
@@
Although growing, the number of professional fact-checkers remains low in comparison to the vast amount of misinformation. The latest census by the Duke Reporters’ Lab [^2] https://reporterslab.org/fact-checking identified 378 active fact-checking projects worldwide. Only 88 organizations worldwide (34 in Europe) actively cooperate with social media platforms through the International Fact-checking network [^3] https://ifcncodeofprinciples.poynter.org/signatories, and in some cases, there is merely one fact-checker per country. Imbalance between such scarcity of fact-checkers and misinformation overload causes just a fraction of potential false content is being checked. Furthermore, the work of fact-checkers is laborious and sometimes repetitive. However, there is an opportunity to improve balance by providing appropriate technological support for challenging or repetitive tasks.
@@
While there is an increasing body of research on automation of the fact-checking process, especially with the utilization of Artificial Intelligence (AI) technologies, full automation of the whole fact-checking process is still viewed with skepticism by the fact-checkers as well as researchers (noa, 2020; Micallef et al., 2022; Dierickx, 2022; Juneja and Mitra, 2022). The main reason is that some parts of the fact-checking process require human judgments or actions. Tasks such as assessing the credibility of the sources of evidence are currently not fully automated with sufficient trustworthiness (e.g., by providing appropriate explanations), accuracy, or generability (e.g., checking complex recent claims where evidence may still be missing) (Nakov et al., 2021a). The real potential of AI-based tools, therefore, currently lies in assisted fact-checking instead of trying to automate the whole process (Graves et al., 2016).
@@
With this in mind, we reinvoke a philosophy of machine design called “autonomation”, developed in the Toyota Production System (Ohno and Bodek, 1988). Autonomation is a blend of “autonomous” and “automation” and may be described as “intelligent automation” or “automation with a human touch”. Compared to (semi-)automation that is typically orchestrated by a centralized computer controller, autonomation separates the work of humans from machines and thus enables the quick correction of the mistakes made by machines. Yet, autonomation relieves humans of the need to continuously judge whether the operation of the machine is right. In the “autonomated” systems, the workers are self-inspecting their work and can source-inspect the work of the machines. This is a difference from an “automatic” system (that operates without human intervention) or an “automated” system (that requires human input or monitoring but is controlled by technology). Autonomation liberates people from automatable tasks, whereas technology has to serve people and processes, not vice versa. In the context of fact-checking, we use the term metaphorically, as it is more commonly used in the engineering industry.
@@
Although the HCI and information science community has significantly helped to better understand the gap between the social and technical, the existing research works on AI-assisted fact-checking are still often detached from real fact-checkers and thus do not optimally comply with their actual needs and expectations. As Nakov et al. point out, there is a “lack of collaboration between researchers and practitioners in defining tasks and developing datasets” for automated fact-checking (Nakov et al., 2021a). This contributes to the skepticism of fact-checkers towards automating the entire fact-checking process, especially processes that require human judgments (noa, 2020; Micallef et al., 2022). On the contrary, some studies, e.g., (Nguyen et al., 2018) mention too much trust in the system’s predictions.
@@
In this study, our objective is to investigate the fact-checking procedures to design the effective and appropriate human-centered AI tools to counter false information. The role of humans is central here: the tools should empower humans instead of replacing them. This work offers the following contributions:
@@
    We unify the various categorizations of the fact-checking process across different computer and social science literature (noa, 2020; Nakov et al., 2021a; Micallef et al., 2022; Barrón-Cedeño et al., 2020).
@@
    We investigate the activities, problems, resources, and tools of under-researched European fact-checkers and examine the specifics of fact-checking work in the non-English speaking-regions. We align them with the knowledge from existing studies, namely (noa, 2020; Nakov et al., 2021a; Dierickx, 2022; Micallef et al., 2022), and visualize them jointly as conceptual models.
@@
    We identify new opportunities for research and development of AI tools designed to support fact-checkers’ daily work in line with human-centered AI principles. In this direction, we discuss three AI tasks that previous studies of fact-checkers’ needs and activities have not addressed yet.
2. Background and related work
2.1. Practices and challenges of fact-checkers in Europe and beyond
Although the research on fact-checkers’ practices and problems is growing (see, e.g.,  (noa, 2020; Micallef et al., 2022; Koliska and Roberts, 2024; Westlund et al., 2024)), the European region as well as non-American and non-English speaking fact-checkers remain still under-explored as can be seen on the map (Fig. 1). At the same time, there are specific challenges that the European fact-checkers face.
These stem from two main sources. First, the European fact-checkers need to operate in a highly multinational and multilingual environment, since many disinformation campaigns easily spread across the borders. To maximize the negative impact in each affected country, disinformation narratives and content itself is often translated and adjusted to specific societal and cultural aspects.
Second, the problem of online disinformation and foreign information manipulations and inference has been recognized in official policy documents at the level of the European Union (EU) and specific legislation has been drafted impacting how the social media platforms should respond to disinformation. Specifically, a European approach to tackle these phenomena emphasizes the importance of a transparent, trustworthy, and accountable online ecosystem while promoting quality journalism and media literacy [^4] https://eur-lex.europa.eu/legal-content/EN/TXT/?uri=celex:52018DC0236. Fact-checking has been recognized as one of the important ways to tackle false information and it was supported by various EU initiatives such as the Social Observatory for Disinformation and Social Media Analysis (SOMA) [^5] https://www.disinfobservatory.org/ or the European Digital Media Observatory (EDMO) [^6] https://edmo.eu/ together with its regional hubs. Most recently, with the adoption of the Digital Service Act (DSA) [^7] https://commission.europa.eu/strategy-and-policy/priorities-2019-2024/europe-fit-digital-age/digital-services-act_en, the previously voluntary Code of Practice on Disinformation [^8] https://digital-strategy.ec.europa.eu/en/library/2022-strengthened-code-practice-disinformation has become the official Code of Conduct, which will be used to determine DSA compliance of the online platforms operating in the EU digital space regarding disinformation risks [^9] https://digital-strategy.ec.europa.eu/en/library/code-conduct-disinformation.
@@ 2.1. Practices and challenges of fact-checkers in Europe and beyond
Figure 1. The regions covered in existing research on fact-checkers (when specified)
@@
Regarding the existing studies on fact-checkers’ practices and problems, the Full Fact report (noa, 2020) was the first one (published in 2020) to globally examine them. Following the semi-structured interviews with fact-checkers from 19 organizations, the report lays out the main challenges fact-checkers face, such as large amounts of potential claims to check, managing tips and suggestions from readers, or the accessibility of information from relevant authorities to check information credibility.
Another study published in 2022 consisted of semi-structured interviews conducted with 21 fact-checkers from 19 countries (Micallef et al., 2022). The results highlight the fact-checkers’ motivations, which are driven by a sense of social responsibility, despite encountering numerous challenges throughout the process. This was further reinforced in (Koliska and Roberts, 2024), which confirms that fact-checkers commonly believe in their ability to determine objective truth, relying on evidence and a transparent process that ensures reproducibility.
Regarding the most common challenges of fact-checkers, these include – based on (Micallef et al., 2022) – the enormous volume of misinformation, the shortcomings in available tools (such as algorithms that favor virality), confirmation bias of information consumers, the existence of echo chambers in online spaces, and insufficient data and resources for fact-checking efforts. Besides that, the authors point out the reasons for the limited uptake of computational tools by these professionals, such as the limited scope to a specific function within the fact-checking process. Furthermore, the integration of the tools with each other proved to be lacking, as the tools are developed by different entities. The authors suggest developing a unified platform with humans-in-the-loop, or, alternatively, establishing a set of standards to turn the fact-checking process into an efficient and streamlined operation. Five key problematic areas regarding misinformation fact-checking were also identified in (Westlund et al., 2024) to be: 1) the limited affordances of digital technologies, 2) limited agency on platform infrastructures, 3) limited expertise and human resources, 4) hostility toward fact-checking actors, and 5) fact-checks fueling misinformation.
Additional studies focused on broadening the understanding of various stakeholders (editors, external fact-checkers, in-house fact-checkers, investigators, and researchers, as well as social media managers and advocates) (Juneja and Mitra, 2022) or on the identification of common features of fact-checking methodology (Sánchez Gonzales et al., 2024).
However, none of these studies focused specifically on the European context. The notable exception is the research on fact-checkers’ practices done by NORDIS (NORdic observatory for digital media and information DISorders) (Dierickx, 2022), which was, however, limited in scope to Nordic countries, i.e., Norway, Sweden, Finland, and Denmark. Semi-structured interviews with 14 respondents from these countries, supplemented with 5 non-Nordic professional journalists and fact-checkers, were conducted. The findings revealed four types of tools that are needed to automate the parts of the fact-checking process considered “boring” for human fact-checkers (social network monitoring, political debate monitoring, claim collection and detection, and context-dependent verification, especially multimedia content) as well as their expected characteristics (e.g., adaptation of tools to Nordic languages). The authors underline the necessity to incorporate the context into the designing process of AI tools, especially 1) the ethical principles of journalism; 2) human values (such as social responsibility to deliver accurate information, creativity, and intuitiveness of the fact-checking process); 3) human expertise with technology; and 4) human relationship to the information resources. In the follow-up study (Dierickx et al., 2023), the authors suggest that the focus needs to be moved from a technological point of view toward a social one, provided that a confidence relationship is established between the communities, developers, and fact-checkers/journalists involved from either side of the tool.
In our study, we increase the scope to organizations from 20 European countries and investigate the differences as well as similarities in their problems and needs across various dimensions (including geography, language, or organizational type).
2.2. AI support of fact-checkers
A lot of AI research is devoted to assisting fact-checkers by automating the individual stages of fact-checkers’ work (Nakov et al., 2021a; Zeng et al., 2021). Reflecting (at least partially) the common steps of the fact-checking process identified in studies above, it focuses on tasks such as finding claims worth fact-checking (Gencheva et al., 2017; Berendt et al., 2021; Konstantinovskiy et al., 2021), searching for previously fact-checked claims (Shaar et al., 2021a; Vo and Lee, 2020; Sheng et al., 2021; Kazemi et al., 2021; Hardalov et al., 2022), or claim verification and evidence retrieval (Zhao et al., 2020; Lee et al., 2020; Popat et al., 2017). Another group of research focuses on the development of the necessary datasets (Shaar et al., 2020; Nakov et al., 2021b; Srba et al., 2022; Hardalov et al., 2022) and end-to-end systems or monitoring platforms (Hassan et al., 2017; Srba, [n. d.]). Developing AI-based methods has also been addressed through competitions and data challenges, especially as a part of CheckThat! Labs and SemEval workshops, where tasks such as check-worthiness estimation (Shaar et al., 2021b) or detecting previously fact-checked claims (Shaar et al., 2021a; Peng et al., 2025) have been proposed.
In parallel with research and development of various AI methods and models, several tools aiming to support the fact-checking process have been created. Some of them directly apply AI in order to provide fact-checkers with more advanced features, including automation of some of the fact-checking steps. In the following overview (see also Fig. 13), we mention especially those that have been explicitly mentioned by the fact-checkers during our study.
At first, there are several tools that allow social media monitoring. Typically, these tools are developed for purposes different from fact-checking (e.g., marketing), such as NewsWhip [^10] https://www.newswhip.com/, StoryBoard [^11] https://storyboard.news/ or BuzzSumo [^12] https://buzzsumo.com/. Fact-checkers, however, found their way to use these tools also for fact-checking purposes. Some tools are also provided directly by social media platforms, such as Meta Content Library [^13] https://transparency.meta.com/en-gb/researchtools/meta-content-library/ (a replacement for the former and widely used CrowdTangle tool) dedicated to monitoring Facebook, Instagram, and Threads, or X Pro [^14] https://pro.x.com/ (a successor of the former TweetDeck tool) dedicated to monitoring X.
For searching within existing fact-checks, fact-checkers can also use dedicated tools. To this end, Google Fact Check Explorer tool [^15] https://toolbox.google.com/factcheck/explorer allows to search within fact-checking articles that use the so-called ClaimReview Schema [^16] https://schema.org/ClaimReview. The ClaimReview Schema allows fact-checkers to annotate their fact-checks by the most important metadata, such as the fact-checked claim, the resulting veracity label, or the link to fact-checked content.
Finally, there are tools that aim to perform end-to-end fact-checking (automating all its steps) or at least cover multiple steps at the same time. At first, Full Fact AI [^17] https://fullfact.org/ai/about/ provides a set of AI-based tools for collecting and monitoring the data, identifying and labeling claims, matching identified claims against a database of past fact-checks, and evidence retrieval to speed up verification of new so-far non-fact-checked claims. Similarly, ClaimBuster [^18] https://idir.uta.edu/claimbuster/ provides support from monitoring and selecting claims up to their verification against the selected knowledge-bases (Hassan et al., 2017). Lastly, the Verification plugin [^19] https://www.veraai.eu/posts/verification-plugin developed and maintained by the EU-funded vera.ai project (while originally created in the InVID and WeVerify projects), provides a set of tools for detecting various types of false information not only in textual content but also in images and videos.
Despite existing support and analyses of fact-checkers’ work, there is a partial disconnection between the fact-checkers’ needs and the efforts to create (AI) tools to support their work. Their design is treated as a technical solution to a technological problem and ignores the complex social and situational context (Juneja and Mitra, 2022). There is also a lack of collaboration between AI researchers and developers with fact-checkers (Nakov et al., 2021a). For example, capabilities of many of the existing fact-checking tools may be limited because various AI-based tools (e.g., detection methods, search engines) typically achieve a worse performance when applied on non-English or low-resource languages. Similarly, the stage monitoring the online space is insufficiently covered by AI research in the context of fact-checkers, although it is carried out routinely by these professionals and in some works (noa, 2020; Juneja and Mitra, 2022) even recognized as “the hardest part of the fact-checking process”. A notable exception is the recent study (Liu et al., 2024), which indicates a change in trends by using a co-design approach to develop NLP tools for fact-checkers.